Determine the length of an aluminum wires that has a resistance of 158 milliohms and a diameter of 0.113. Resistivity of aluminum = 2.83x10^-8


l=


1
Expert's answer
2021-02-08T18:37:40-0500

The length of the wire can be defined as follows:


"R=\dfrac{\rho L}{A},""L=\dfrac{RA}{\rho}=\dfrac{158\cdot10^{-3}\ \Omega\cdot \dfrac{\pi\cdot(1.13\cdot10^{-4}\ m)^2}{4}}{2.83\cdot10^{-8}\ \Omega\cdot m}=0.056\ m."